OpenStudy (anonymous):

Solve the following system of equations. 3x - 2y = 6 6x - 4y = 12 (0, 0) (6, 12) Infinitely Many Solutions No Solutions

OpenStudy (anonymous):

is it the 3rd one?

OpenStudy (anonymous):

you need to substitute the values for e.g. 3x-2y=6-------eq1 6x-4y=12------eq2 now take eq1 and find the value of x 3x-2y=6 3x=6+2y x=6+2y/3 now substitute the value of x in eq2 this way you will be able to find the exact value of y then substitute this vakue of y in eq1 to find exact the value of x can you do it now?

OpenStudy (anonymous):

is it the 3rd one?

OpenStudy (anonymous):

have you solved it the way i told you?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so you are getting the third option?

OpenStudy (anonymous):

thats what i came out with

OpenStudy (anonymous):

yes its right